(Appendix 4B) Direct Method of Support Department Cost Allocation

Stevenson Company is divided into two operating divisions: Battery and Small Motors. The company allocates power and general factory costs to each operating division using the direct method. Power costs are allocated on the basis of the number of machine hours and general factory costs on the basis of square footage. Support department cost allocations using the direct method are based on the following data:

Support Departments Operating Divisions

Power
General Factory
Battery
Small Motors
Overhead costs $160,000 $430,000 $163,000 $84,600
Machine hours 2,000 2,000 8,000 2,000
Square footage 1,000 1,500 5,000 17,500
Direct labor hours 18,000 60,000

Required:

1. Calculate the allocation ratios for Power and General Factory. (Note: Carry these calculations out to four decimal places, if necessary.)

Power General Factory
Battery
Small Motors

2. Allocate the support service costs to the operating divisions. Round all amounts to the nearest dollar. Use a minus sign to indicate a subtraction. For those boxes in which no entry is required, leave the box blank or enter zero ("0").

Support Departments Operating Divisions
Power General Factory Battery Small Motors
Direct costs $ $ $ $
Allocate:
    Power
    General Factory
Total $ $ $ $

3. Assume divisional overhead rates are based on direct labor hours. Calculate the overhead rate for the Battery Division and for the Small Motors Division. Round your answers to the nearest cent.

Battery overhead rate $
Small Motors overhead rate $
Answer #1

Answer

  • Requirement 1

Power [8000 + 2000 = 10000 machine hours]

General factory [5000 + 17500 = 22500 sq feet]

Battery

0.8000

0.2222

Small Motors

0.2000

0.7778

--Workings

Power [8000 + 2000 = 10000 machine hours]

General factory [5000 + 17500 = 22500 sq feet]

Battery

=8000/10000

=5000/22500

Small Motors

=2000/10000

=17500/22500

  • Requirement 2

Power

General Factory

Battery

Small Motors

Direct Cost

$160,000

$430,000

$163,000

$84,600

Allocate:

Power

($160,000)

$0

$128,000

$32,000

General factory

($430,000)

$95,546

$334,454

Total

$0

$0

$386,546

$451,054

--Working

Power

General Factory

Battery

Small Motors

Direct Cost

160000

430000

163000

84600

Allocate:

Power

-160000

0

=160000*0.8

=160000*0.2

General factory

-430000

=430000*0.2222

=430000*0.7778

  • Requirement 3

Total Overhead after allocation

Tota Direct Labor hours

Overhead rates = Answer

Battery

$386,546

                                     18,000

$                                21.47

Small Motors

$451,054

                                     60,000

$                                  7.52